OpenStudy (anonymous):

An attractive force of 7.2 N occurs between two point charges that are 0.05 m apart. If one charge is - 4.0 µC, what is the other charge? A. +0.5 µC B. +2.0 µC C. -2.0 µC D. -4.0 µC

OpenStudy (anonymous):

@Michele_Laino :)

OpenStudy (michele_laino):

here we have to apply the law of Coulomb: \[\Large F = K\frac{{{q_1}{q_2}}}{{{d^2}}}\]

OpenStudy (anonymous):

ok! what do we plug in?

rvc (rvc):

do you know for what the symbols stand @iheartfood

OpenStudy (anonymous):

yes, i believe so.... i just get confused with inserting :/

rvc (rvc):

oh tell us what the symbols mean

OpenStudy (michele_laino):

the next step is: \[\large {q_2} = \frac{{F{d^2}}}{{K{q_1}}} = \frac{{7.2 \times {{\left( {5 \times {{10}^{ - 2}}} \right)}^2}}}{{9 \times {{10}^9} \times 4 \times {{10}^{ - 6}}}} = ...Coulombs\]

OpenStudy (anonymous):

we get 5E-7?

OpenStudy (anonymous):

is that right? if so, what is the next step? :)

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

what is the sign of q_2?

OpenStudy (anonymous):

positive?

OpenStudy (michele_laino):

yes! since the force is attractive and q_1 is negative

rvc (rvc):

attractive force exists between opposite charrges

OpenStudy (anonymous):

ooh yay! so our solution will be either -2.0 or -4.0 ? how can we find which it is?

OpenStudy (michele_laino):

our solution is: \[\large 5 \times {10^{ - 7}} = 0.5 \times {10^{ - 6}} = 0.5\mu C\]

rvc (rvc):

no if one is negative other will be positive

OpenStudy (anonymous):

oh ok! yay thank you!! +0.5 it is!!

OpenStudy (michele_laino):

yes! @rvc

rvc (rvc):

thanks

OpenStudy (michele_laino):

that's right! @iheartfood